OpenStudy (anonymous):

Show the work to solve |3x - 2| < 7 Describe the graph of the solution in words

OpenStudy (anonymous):

just like last one

OpenStudy (anonymous):

one compound inequality \[-7<3x-2<7\] add 2 and divide by 3 to get your answer

OpenStudy (anonymous):

-7 < 3x - 2 < 7 -5 < 3x < 9 -5/3 < x < 3\[x \in (-5/3, 3)\]

OpenStudy (anonymous):

\[-5<3x<9\] then what abtrehearn said

OpenStudy (anonymous):

How do i describe that in words?

OpenStudy (anonymous):

the graph*

OpenStudy (anonymous):

The solution set is all real numbers between -5/2 and 3.

OpenStudy (anonymous):

The graph on the real number line is a line segment starting at -5/3 and ending at 3, and its ends are closed off by parentheses to denote exclusion of endpoints from solution set.

OpenStudy (anonymous):

Can you help with this? Give a unique example of each: 1. An absolute value equation with two solutions 2. An absolute value equation with one solution 3. An absolute value equation with no solutions

OpenStudy (anonymous):

1. 2|3x + 8| + 3 = 7 2, |5x + 6| = 0 3. |71 x + 17| = -29